# EdgeAlpha: Bringing Process Discovery to the Data Sources

### v1

@@ 4.2 EdgeAlpha Algorithm
This phase begins when an outside entity, for example a central node, requests a process model. During this phase, we request the local FMs of all nodes and start or end events within any case, see Figure 3. Subsequently, each node responds with its local FM. Before sending, the node changes all entries that are greater than 0 to 1, which aggregates the data even more and pushes it back into the conventional FM structure. To determine whether its respective activity is an end activity, it verifies whether it detected an event lacking a successor. The requesting unit receives all responses, it assembles the footprint matrix $FM^{all}\in\{1,0\}^{n\times n}$ concatenating all columns $FM^{i}$, $i\in\{0,\dots,n-1\}$ of all $n$ nodes, ordered by activity identifier. Further, we merge the start and end activities into the sets $T_{I}$ and $T_{O}$, respectively, following the conventions of the Alpha Miner. Subsequently, the requesting unit proceeds with the remainder of the algorithm, starting from step four, the determination of $X_{L}$, which includes calculating the FM, here $FM^{all}$, see Algorithm 1.
@@
Figure 3: Phase 2 – Requesting Process Data: First, we request the partial FMs and whether the activity was the start or end of any trace from all nodes. Upon receiving the information, we concatenate the matrices, form start and end activity sets, compute the process model.
4.2.3 Most-Frequent-Predecessors (MFPs).
@@
In phase 1 of EdgeAlpha, the event detecting node sends out requests to all other nodes, resulting in $n-1$ requests where $n$ denotes the number of distinct nodes, i.e., activities in the event log. To reduce communication overhead, we search for the predecessor as follows: we query nodes in order of their likeliness to be the predecessor. Thus, when searching for the predecessor of an event, a node first contacts the most likely predecessor and then the second most likely one and so on. For this, each node records how often each predecessor occurs. Given that activities are, in real-world scenarios, preceded by a consistent subset of activities, this approach prioritizes querying the most frequently occurring predecessors. When detecting a start event, the corresponding node has to still query every single node in the network, to be sure that it is the start event. In our evaluation in Section 5 we show, for real-world event logs, that our statistical approach significantly reduces communication cost and thereby ensures the scalability of EdgeAlpha.
4.2.4 Batching.
@@
Claim
Merging the partial footprint matrices $FM^{0},\dots,FM^{n-1}$ by concatenating them ordered by activity identifier results in the same FM as the $FM^{all}$, formally $\begin{pmatrix}FM^{0}&~{}\dots~{}&FM^{n-1}\end{pmatrix}~{}=~{}FM^{all}.$
Proof
To prove the claim, we show that all direct successions represented in the partial FMs can be found in $FM^{all}$ and vice versa. Thus, proving the claim of equivalence between the merged partial FMs and the FM with the unrestricted knowledge.
@@
Let $i>_{L_{j}}j,~{}i,j\in\{0,\dots,n-1\}$ be a direct succession encoded in the merged partial footprint matrix $\begin{pmatrix}FM^{0}&~{}\dots~{}&FM^{n-1}\end{pmatrix}$. Due to the concatenation of the partial matrices, it holds $FM^{j}_{i}=1$. $FM^{j}$ contains all direct successions between its own activity and the activities preceding it. This means, there is a trace in the event log $L$ in which $i$ is followed by $j$. By definition, $FM^{all}$ captures all direct successions in $L$, meaning it must hold $FM^{all}_{ij}=1$.
@@
It remains to show that the opposite direction holds. Let $k>_{L}\ell,~{}k,\ell\in\{0,\dots,n-1\}$ be a direct succession represented in $FM^{all}$, formally $FM^{all}_{k\ell}=1$. Since each node keeps track of all its predecessor activities in its local FM, we know that $FM^{\ell}_{k}=1$. Merging all partial FMs only means concatenating them ordered by activity identifier. Therefore, it follows $\begin{pmatrix}FM^{0}&~{}\dots~{}&FM^{n-1}\end{pmatrix}_{k\ell}=1$. Thus, proving the claim.
$\Box$
4.2.6 Design Discussion.
@@
A major advantage of EdgeAlpha is that event logs are no longer stored in one central place. Instead, in EdgeAlpha, each node stores its local view on the process, i.e, its partial process information, and only sends aggregates (partial FMs) to a central entity. In addition, we use batching, i.e. a message is not sent after every event, but after a certain time or number of events in order to make exact times and case sequences unrecognizable.
@@
Further, the design of EdgeAlpha is directly applicable to process mining algorithms beyond the Alpha Miner. Algorithms that start by building a FM or keep track of direct successions in the event log, can be projected onto our distribution approach. A well-known example of another algorithm is the Heuristics Miner [22]. The vast majority of results that we show in our evaluation in Section 5, such as the efficiency of Phase 1 of EdgeAlpha, is not specific to the Alpha Miner and directly applies, for example, to the Heuristics Miner.
5 Evaluation
@@
In this section, we evaluate EdgeAlpha in terms of performance. We start by introducing our implementation in Section 5.1 and suitable baselines in Section 5.2. In Section 5.3, we introduce on which well-known datasets we conduct the experiments. We analyze these datasets in Section 5.4. In Section 5.5, we examine how MFP Requesting reduces the number of queried nodes for Phase 1 of the algorithm compared to naive querying of all nodes. We continue with an estimation on what FM sizes to expect during an algorithm run in Section 5.6. We finish the section by discussing the overall scalability of EdgeAlpha in Section 5.7
5.1 Implementation
We implement EdgeAlpha in Python and use a network of docker containers to simulate the distributed nature of EdgeAlpha. We also validate that the resulting process model of EdgeAlpha is identical to the one computed by pm4py [^1] https://pm4py.fit.fraunhofer.de using both artificial datasets and common real-world event logs, such as the datasets in Section 5.3.
5.2 Baselines
@@
In our experiments, we compare EdgeAlpha with and without MFP Requesting. Not using MFP Requesting, i.e. querying all nodes (”Query All”), serves as upper bound. On the lower end, a theoretical, oracle-like, approach bounds the number of queried nodes in EdgeAlpha. It requires no requests if the predecessor corresponds to the node itself and otherwise takes exactly one request to identify the predecessor.
@@
EdgeAlpha assumes a setting in which event detection is part of the algorithm, whereas existing algorithms assume the availability of an event log from the beginning. Thus, comparing EdgeAlpha to the traditional Alpha Miner [4] or other distributed Alpha Miners such as the Map-Reduce approach [11] or the Alpha Miner in the healthcare scenario [14] is not viable. The novelty of EdgeAlpha lies in leveraging the event detection phase to pre-process events, a feature absent in other approaches.
5.3 Datasets
@@
We conduct experiments on the datasets Hospital Log [9], Sepsis Cases [18] and an IoT-Enriched Event Log for Process Mining in Smart Factories [^2] We only consider the MainProcess.xes file of the not cleaned version. [17]. These datasets naturally support geographical distribution over multiple nodes. To further substantiate our evaluation, we also include the BPI Challenge 2017 [10] and the Road Traffic Fine Management Process [16]. Table 1 shows general statistics of the datasets, including the number of events, the number of cases, the average case length, and the number of activities. As each node is responsible for a single activity, the dataset, for example, represents IoT networks with size of 11 nodes in case of the Traffic Fine dataset and 624 nodes in case of the Hospital Log.
@@
Traffic Fine 561,470 150,370 3.73 11
Sepsis 15,214 1,050 14.49 16
Smart Factories 8,607 271 31.76 21
BPIC 2017 1,202,267 31,509 38.16 26
Hospital 150,291 1,143 131.49 624
Table 1: Key data of the datasets showcasing their basic properties such as event counts, the number of cases, the average case length or the number of activities.
Event Log ##\##Start Activities ##\##Self-Loops Avg ##\##pred. (SD) Avg MFP Ratio (SD) Reduction
Traffic Fine 1 4,306 6.36 (2.71) 0.80 (0.17) 64.94%
Sepsis 6 1,034 7.19 (2.67) 0.53 (0.20) 81.30%
Smart Factories 1 5,854 4.62 (3.33) 0.69 (0.07) 92.15%
BPIC 2017 1 448,205 6.85 (5.30) 0.79 (0.18) 93.40%
Hospital 29 26,542 6.78 (12.25) 0.69 (0.28) 96.13%
@@
Table 2: Dataset details that directly influence the number of requests in phase 1 of EdgeAlpha. These include the number of start activities and self-loops as well as details about the predecessors, such as the average number of predecessors or the average ratio of the most frequent predecessor per activity. Additionally, it shows the reduction EdgeAlpha achieves in the average number of queried nodes when using MFP Requesting compared to querying all nodes in the network.
5.4 Dataset Analysis
We first analyze the properties of real-world event logs, see Table 2. For instance, our analysis shows that the Smart Factories event log contains numerous self-loops, which arise due to events having different lifecycle transitions and states in this event log. This characteristic is essential for EdgeAlpha, as self-loops do not require communication. Further, the table shows that activities commonly have only a few predecessors, underlining the potential of MFP Requesting.
5.5 Number of Queried Nodes
@@
To assess EdgeAlpha’s performance, we analyze its two phases separately. The event ordering phase depends highly on the number of potential predecessors a node needs to query. In datasets like the Hospital Log there are up to 623 nodes to query, the network traffic can be considerably large, particularly during concurrent events. MFP Requesting in EdgeAlpha builds on past event data to optimize request in the future. For the first experiments we assume a batch size of one to focus on other characteristics. As we show in the the batch size analysis, the different batch sizes strongly influence the number of queried nodes.
@@
Figure 4: Comparison of average queried nodes with MFP Requesting compared to querying all nodes. EdgeAlpha is able to reduce communication by querying the most frequent predecessors. With knowledge of the start events, EdgeAlpha can save even more communication bandwidth.
@@
Figure 5: Cumulative Distribution Functions illustrating the relationship between the normalized number of requested nodes and the probability of occurrence. This diagram shows the distribution of queried node counts. A steeper curve, such as for the Hospital Log, indicates a higher event proportion necessitating only a small fraction of queried nodes to identify their predecessors, reflecting greater efficiency.
@@
Figure 6: Relationship between batch size and average number of queried nodes per event across all datasets. To be able to compare the datasets with each other, we normalize the average number of queried nodes. Across all datasets, transitioning from a batch size of 1 to 10 results in a more significant reduction in queries than from 40 to 50. With increasing batch size the reduction rate decreases.
@@
Figure 7: Stabilization of queried node count over time in Hospital Log. Since the data is fluctuating, we use moving averages to visualize the stabilization. A reference line (Query All) indicates the maximum number of requested nodes per event. The moving average stabilizes after fewer than 3,000 events. We plot an additional line with a larger window size to show the trend of the stabilization. The plots are limited to the first 6,000 events.
@@
Figure 8: Bar diagram based on the Sepsis Cases illustrating the average number of queried nodes per activity (left y-axis) alongside the number of occurrences of each activity (right y-axis). The x-axis displays the different activities of the event log. To put the averages into relation, we plot the number of requested nodes when querying all nodes as well as the average number of queried nodes over all events. The diagram indicates that the events that query on average the most nodes, occur the least. Start activities are excluded from this conclusion.
5.5.1 Comparison Across Datasets.
@@
Figure 4 illustrates, for each dataset, the average number of nodes queried per event to find the predecessor. Query All denotes the naive way to message every node in the network. This serves as baseline to compare to. With MFP Requesting, nodes require significantly fewer requests to identify their predecessor event. We achieve further message reduction when the process has a designated start event, making predecessor requests for the respective node redundant. Table 2 also summarizes the significant reduction in the number of queried nodes. Notably, this reduction increases with the number of activities, influenced by factors such as an expanding average case length, diminishing significance of start events, varying numbers of self-loops relative to total events, and the presence of just a few highly frequent predecessors.
@@
Figure 6 presents Cumulative Distribution Functions (CDF) comparing all datasets. The Hospital Log shows the steepest increase with over 90% of events requiring queries to no more than 5% of the nodes. Conversely, the Road Traffic Fine Management Process has barely any occurrences where the predecessor can be identified after querying just 5% of the nodes. This deviation stems from the limited number of 11 activities in the Traffic Fine dataset. Hence, within the 5% threshold, only self-loops are considered.
5.5.2 Stabilization Analysis.

### v3

@@ IV-E Phase 2 – Requesting Process Data
This phase begins when an outside entity, for example, a central node, requests a process model. During this phase, we request the local FMs of all nodes and start or end events within any case, see Fig. 3. Subsequently, each node responds with its local FM. Before sending, the node changes all entries greater than 0 to 1, aggregating the data even more. This converts the local partial FM back into the conventional FM structure. Each node checks if events without a successor were recorded to determine whether its respective activity is an end activity. The requesting entity receives all responses, and it assembles the footprint matrix $FM^{all}\in\{1,0\}^{n\times n}$ concatenating the columns $FM^{i}$, $i\in\{0,\dots,n-1\}$ of all $n$ nodes, ordered by activity ID. Further, we merge the start and end activities into two sets $T_{I}$ and $T_{O}$, following the conventions of the Alpha Miner. Next, the requesting unit proceeds with the remainder of the Alpha Miner algorithm, starting after the calculation of the FM, here $FM^{all}$.
@@
Figure 3: Phase 2 – Requesting Process Data: We request partial FMs and start/end activity flags from all nodes. Upon receiving the data, we concatenate the matrices, form start and end activity sets, and compute the process model.
IV-F Most-Frequent-Predecessors (MFPs)
@@
In Phase 1 of EdgeAlpha, the event detecting node sends out requests to all other nodes, resulting in $n-1$ requests where $n$ denotes the number of distinct nodes, i.e., activities in the event log. To reduce communication overhead, we search for the predecessor by querying nodes in order of their likelihood of being the predecessor. Thus, when searching for an event’s predecessor, a node first contacts the most likely predecessor, then the second most likely one, and so on. For this, during runtime each node continuously records how often each predecessor occurs. As we show later in a dataset analysis, most activities in real-world scenarios are preceded by a typically stable and small subset of activities, this approach prioritizes querying the most frequently occurring predecessors. However, when detecting a start event, the corresponding node still has to query every node in the network to ensure it is the start event. In our evaluation in Section V we show, for real-world event logs, that our statistical approach significantly reduces communication cost and thereby ensures scalability.
IV-G Batching
@@ IV-H Correctness
Claim 1
Merging the partial footprint matrices $FM^{0},\dots,FM^{n-1}$ by concatenating them ordered by activity ID results in the same FM as the $FM^{all}$, formally $\begin{pmatrix}FM^{0}&~{}\dots~{}&FM^{n-1}\end{pmatrix}~{}=~{}FM^{all}.$
Proof 1
To prove the claim, we show that all direct successions represented in the partial FMs can be found in $FM^{all}$ and vice versa. Thus proving the claim of equivalence between the merged partial FMs and the FM with unrestricted knowledge.
@@
Let $i>_{L_{j}}j,~{}i,j\in\{0,\dots,n-1\}$ be a direct succession encoded in the merged partial footprint matrix $\begin{pmatrix}FM^{0}&~{}\dots~{}&FM^{n-1}\end{pmatrix}$. Due to the concatenation of the partial matrices, it holds $FM^{j}_{i}=1$. $FM^{j}$ contains all direct successions between its activity and its preceding activities. This means there is a trace in the event log $L$ in which $i$ is followed by $j$. By definition, $FM^{all}$ captures all direct successions in $L$, meaning it must hold $FM^{all}_{ij}=1$.
@@
It remains to show that the opposite direction holds. Let $k>_{L}\ell,~{}k,\ell\in\{0,\dots,n-1\}$ be a direct succession represented in $FM^{all}$, formally $FM^{all}_{k\ell}=1$. Since each node keeps track of all its predecessor activities in its local FM, we know that $FM^{\ell}_{k}=1$. Merging all partial FMs only means concatenating them ordered by activity ID. Therefore, it follows $\begin{pmatrix}FM^{0}&~{}\dots~{}&FM^{n-1}\end{pmatrix}_{k\ell}=1$. Thus proving the claim. $\Box$
IV-I Design Discussion
@@
A major advantage of EdgeAlpha is that event logs are no longer stored in one central place. Instead each node stores its local view on the process, i.e, its partial process information, and only sends aggregates (partial FMs) to a central entity. In addition, we use batching, i.e. a message is not sent after every event, but after a certain time or number of events in order to make exact times and case sequences unrecognizable.
@@
Further, the design of EdgeAlpha directly applies to process mining algorithms beyond the Alpha Miner. Algorithms that start by building an FM or keeping track of direct successions in the event log can be projected onto our distribution approach. A well-known example of another algorithm is the Heuristics Miner [14]. The vast majority of results that we show in our evaluation in Section V, such as the efficiency of Phase 1 of EdgeAlpha, are not specific to the Alpha Miner and directly apply, for example, to the Heuristics Miner.
V Evaluation
@@
In this section, we evaluate EdgeAlpha in terms of correctness, accuracy, and performance. We present our implementation in Section V-A and suitable baselines in Section V-B. Section V-C introduces on which well-known datasets we conduct the experiments. We analyze these datasets in Section V-D. In Section V-E we discuss correctness experiments and the compliance between the FMs of EdgeAlpha and the traditional Alpha Miner. We examine how MFP Requesting reduces the number of queried nodes for Phase 1 of the algorithm compared to naive querying in Section V-F. Section V-G estimates FM sizes during an algorithm run and Section V-H discusses the overall scalability of EdgeAlpha.
V-A Implementation
We implement EdgeAlpha in Python and use a network of docker containers to simulate the distributed nature of EdgeAlpha.
V-B Baselines
@@
In our experiments, we compare EdgeAlpha with and without MFP Requesting. ”Query All” represents the upper bound where all nodes are queried, while a theoretical oracle-like approach serves as the lower bound. The latter requires no requests if the predecessor corresponds to the node itself, and only one request to identify the predecessor otherwise.
@@
EdgeAlpha’s design assumes event detection as part of the algorithm, whereas existing algorithms assume the availability of an event log from the beginning. Thus, comparing EdgeAlpha to the traditional Alpha Miner [12] or other distributed Alpha Miners such as the Map-Reduce approach [19] or the Alpha Miner in the healthcare scenario [9] is not viable. The novelty of EdgeAlpha lies in utilizing the event detection phase to pre-process events, a feature absent in other approaches.
V-C Datasets
@@
We conduct experiments on the datasets Hospital Log [13], Sepsis Cases [24] and an IoT-Enriched Event Log for Process Mining in Smart Factories [^2] We only consider the MainProcess.xes file of the not cleaned version. [3]. These datasets naturally support geographical distribution over multiple nodes. To further substantiate our evaluation, we also include the BPI Challenge 2017 [25] and the Road Traffic Fine Management Process [26]. In our setting, each activity in a dataset represents a node in an IoT network.
TABLE I: Dataset Characteristics (Standard deviation in parentheses if not stated otherwise).
@@
Traffic Fine 561,470 11 (1) 150,370 3.73 4,306 6.36 (2.71) 0.80 (0.17) 64.94%
Sepsis 15,214 16 (6) 1,050 14.49 1,034 7.19 (2.67) 0.53 (0.20) 81.30%
Smart Factories 8,607 21 (1) 271 31.76 5,854 4.62 (3.33) 0.69 (0.07) 92.15%
BPIC 2017 1,202,267 26 (1) 31,509 38.16 448,205 6.85 (5.30) 0.79 (0.18) 93.40%
Hospital 150,291 624 (29) 1,143 131.49 26,542 6.78 (12.25) 0.69 (0.28) 96.13%
V-D Dataset Analysis
@@
We analyze the properties of real-world event logs. Next to basic information such as the numbers of events, activities and cases and the average case length, Table I includes data that directly influences the number of requests in Phase 1. These include the number of start activities and self-loops as well as details about the predecessors, such as the average number of predecessors or the average ratio of the most frequent predecessor per activity. Additionally, it shows the reduction EdgeAlpha achieves in the average number of queried nodes when using MFP Requesting compared to querying all nodes in the network.
Even in a dataset like the Hospital Log with over 600 distinct activities, the average number of preceding activities per activity stays below 7, showcasing the potential of MFP Requesting to reduce the average number of queried nodes per event. The dataset analysis reveals that event logs like the Smart Factories dataset contain numerous self-loops, resulting from different lifecycle transitions and states within the recorded events. This characteristic benefits EdgeAlpha, as self-loops do not require communication. For self-loops, too, only in case of communication delays a self-loop may be set wrongfully for which, however, EdgeAlpha provides a rectification procedure.
V-E Experimental Correctness and Fitness Analysis
As a first step, we validate that the resulting process model of EdgeAlpha is identical to the one computed by pm4py [^3] https://pm4py.fit.fraunhofer.de using both artificial datasets and common real-world event logs, such as the datasets in Section V-C.
Next, we examine how the merged FMs obtained by querying the nodes for their partial FMs after an increasing number of events differ from the intermediate FM produced by the traditional Alpha Miner. We compare them by using the footprint-based fitness value implemented in pm4py. Fig. 7 confirms the full compliance of EdgeAlpha’s resulting FM with the Alpha Miner’s FM across all dataset. It also demonstrates that nearly all datasets reach 90% fitness with less than 5% of events. Here, the Smart Factories dataset is an exception, reaching 90% fitness after processing approximately 41% of the events, although it is considerably small, as shown in Table I. The results indicate that querying interim results can be highly informative, potentially eliminating the need to process all events there are typically present in a common event log before requesting the first process model. Moreover, our analysis suggests that reducing the sliding window for discarding old data would have minimal impact on the resulting process model, but is dependent on the current dataset.
V-F Number of Queried Nodes
@@
To assess EdgeAlpha’s performance, we analyze its two phases separately. The event ordering phase depends highly on the number of potential predecessors a node queries, leading to significant network traffic, particularly during concurrent events, in datasets like the Hospital Log, where up to 623 nodes are queried. MFP Requesting in EdgeAlpha builds on past event data to optimize requests in the future. For the first experiments, we assume a batch size of one to focus on other characteristics. As our batch size analysis shows, the different batch sizes strongly influence the number of queried nodes.
@@
Figure 4: Comparison of average queried nodes with MFP Requesting and querying all nodes. Diagram also shows standard error. EdgeAlpha can reduce communication by querying the most frequent predecessors. With knowledge of the start events, EdgeAlpha can save even more communication bandwidth.
Figure 5: Fitness over time. Intermediate result FMs of the data sets quickly converge into the FM of the traditional Alpha Miner. The BPIC 2017 data set, for example, already has a fitness value of over 90% after 200 events compared with the Alpha Miner FM, which is approx. 0.02% of the events in the event log.
@@
Figure 6: Cumulative Distribution Functions illustrating the relationship between the normalized number of requested nodes and the probability of occurrence. A steeper curve indicates a higher event proportion necessitating only a small fraction of queried nodes to identify their predecessors, reflecting greater efficiency.
@@
Figure 7: Relationship between batch size and average number of queried nodes per event. Normalizing the average queried nodes for comparison. Across all datasets, transitioning from a batch size of 1 to 10 results in a more significant reduction in queries than from 40 to 50. With increasing batch size, the reduction rate decreases.
@@
Figure 8: Stabilization of queried node count over time in Hospital Log. Since the data fluctuates, we use moving averages to visualize the stabilization. A reference line (Query All) indicates the maximum number of requested nodes per event. The moving average stabilizes after fewer than 3,000 events. We plot an additional line with a larger window size to show the trend of the stabilization. The plots are limited to the first 6,000 events.
@@
Figure 9: Bar diagram based on the Sepsis Cases illustrating the average number of queried nodes per activity (left y-axis) alongside the number of occurrences of each activity (right y-axis). The x-axis displays the different activities of the event log and highlights the start activities (S). We plot the number of requested nodes when querying all nodes and the average number of queried nodes over all event. The diagram shows that activities with the highest average queries occur the least, excluding start activities.
V-F1 Comparison Across Datasets
@@
Fig. 4 illustrates, for each dataset, the average number of nodes queried per event to find the predecessor. Query All denotes the naive way to message every node in the network. This serves as a baseline to compare to. With MFP Requesting, nodes require significantly fewer requests to identify their predecessor event. We achieve further message reduction when the process has a designated start event, making predecessor requests for the respective node redundant. Table I shows the significant reduction in the number of queried nodes. This reduction increases with the number of activities, influenced by factors such as an expanding average case length, diminishing significance of start events, varying numbers of self-loops relative to total events, and the presence of just a few highly frequent predecessors.
@@
Fig. 7 shows Cumulative Distribution Functions (CDF) for all datasets. The Hospital Log has the steepest increase with over 90% of events needing queries to no more than 5% of the nodes. Conversely, the Traffic Fine dataset has few instances where the predecessor can be identified after querying just 5% of the nodes. This deviation stems from the limited number of 11 activities in the Traffic Fine dataset. Hence, within the 5% threshold, only self-loops are considered.
V-F2 Stabilization Analysis